OpenStudy (anonymous):

By using the Ideal gas law variables and by using the variables m to represent mass and M to represent molar mass, derive an algebraic expression for this: the gas density (vapor density) of a gas without using "m" or "V"

OpenStudy (aaronq):

PV=nRT n=m/M PV=mRT/M P=mRT/MV d=m/V P=dRT/M d=PM/RT